A given wire has a resistance of 32 ohms. If its length is 872 m., how much length must be


cut-off from the wire in order to reduce its resistance to 13.4 ohms?

1
Expert's answer
2022-03-21T12:45:03-0400

We know that

Resistance

"R=\frac{\rho l}{A}"

"R_1=\frac{\rho l_1}{A}"

Case (2)

"R_2=\frac{\rho l_2}{A}"

"\frac{R_1}{R_2}=\frac{l_1}{l_2}"

"l_2=\frac{R_2}{R_1}l_1"

"l_2=\frac{13.4}{32}\times872=365.15m"